If cos (α + β) = 0, then sin (α – β) can be reduced to [NCERT Exemplar Problems]
(a) cos β
(b) cos 2β
(c) sin α
(d) sin 2α

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Answer: b Explaination: (b) cos (α + β) = 0 = cos 90° ⇒ α + β = 90° sin (α – β) = sin (α – β + β – β) = sin (α + β – 2β) = sin (90° – 2β) = cos 2β